# Logical perspectives on learning statistical objects

### v1

@@ 5.5 Extension to the real-valued case
$\frac{\log^{2}(an)}{bn}\leq\frac{\left(\log C+\log(a^{2}b^{-1})+\log\log^{2}(ab^{-1})\right)^{2}}{bC(ab^{-1})\log^{2}(ab^{-1})}\leq\frac{\left(\log C+4\log(ab^{-1})\right)^{2}}{aC\log^{2}(ab^{-1})}\leq\frac{(\log C/\log(ab^{-1})+4)^{2}}{C}\leq\frac{(\log C+4)^{2}}{C},$
@@
which completes the proof. ∎
Theorem 5 follows from Theorem 3 using Fact 11.
5.6 Simpler and better VC bounds for the distribution function class
A much simpler argument is available that provides bounds for the distribution class or dual distribution class. We explain the idea for the dual distribution class formed over a concept class. We know that if a concept class is agnostic PAC learnable, then so is its dual class, where the concepts are given by elements $x$ of the range space. We can then conclude by routine calculation with dimensions, that for each $k$, the functions on the parameter space given by normalized sums of elements $\frac{x_{1}+\ldots+x_{k}}{k}$ is agnostic PAC learnable. But by a basic result in statistical learning theory, the fact that our original concept class is PAC learnable means we can approximate each function in the dual distribution class arbitrarily closely by normalized sums.
Recall that in the special case of a concept class, the result for the dual distribution class had been proven in prior work (Hu et al., 2022). The simpler proof we present here, like the analytic proof that goes via the randomization class, improves on the bound given in prior work. We now explain the idea of this alternative approach.
@@ 5.6 Simpler and better VC bounds for the distribution function class
This fits the composition framework in Theorem 1 of (Attias & Kontorovich, 2024).
Fact 18.
Attias & Kontorovich (2024) The $\gamma$ fat-shattering dimension of the composed class $\chi^{\mathcal{C}}_{m}$ is bounded by:
@@
We now show that this approach generalizes easily from concept classes to function classes. That is, we give an alternative proof of the preservation of agnostic PAC learnability, without going through randomization. We do not compute sample complexity bounds explicitly for this alternative proof, but they are similar to those given above.
Theorem 4.
@@
This again fits the composition framework in Theorem 1 of (Attias & Kontorovich, 2024). From the theorem we have the $\gamma$ fat-shattering dimension of the composed class is bounded by:
@@ 5.7 The realizable case for PAC learning
Randomization, and the related distribution class constructions, do not preserve PAC learnability in the realizable case.
Proposition 5.
There is a hypothesis class ${\mathcal{H}}$ that is realizable PAC learnable, but the distribution function class, dual distribution function class, and random variable class based on ${\mathcal{H}}$ are not realizable PAC learnable.
@@
Proof.
A dual class element is given by an $x_{0}$ in the range space, with any other element in the same partition $X_{i}$ as $x_{0}$ inducing the same function. If we see samples $\vec{b}_{1}\ldots\vec{b}_{n}$, with value at most $\frac{1}{8}$, we will be able to exclude some partition elements $X_{i}$, but we will have infinitely many $X_{i}$ possible. ∎
@@
Lemma 7.
The distribution class and (hence) the random variable class of ${\mathcal{H}}_{0}$ are not PAC learnable in the realizable case. In fact, we may simply look at the class on $X$ of “two choice distributions”, consisting of all hypotheses $\lambda\cdot h_{b}+(1-\lambda)\cdot h_{b^{\prime}}$ for rational $\lambda\in[0,1]$ with $b,b^{\prime}\in B$.
Proof.
We prove this by showing that this new class has infinite $\frac{1}{8}$-graph dimension, as defined in (Attias et al., 2023), which we now review. Fix a natural number $n$. For our purposes, we only need to know when a class has $\frac{1}{4}$-graph dimension at least $n$. The definition of graph dimension states that this holds when we have $x_{0},\dots,x_{n-1}\in X$, $f_{1},\dots,f_{n-1}\in[0,1]$, and for each $\beta\in\{0,1\}^{n}$, a hypothesis $h^{\prime}_{\beta}$ such that
@@
Note that, unlike with fat-shattering, we have an equality for the zero values of a branch and an inequality for the one value. Theorem 1 of (Attias et al., 2023) shows that an infinite $\frac{1}{4}$-graph dimension — that is, having such witnesses for each $n$ — implies that the class is not realizable PAC learnable. To find the required witnesses, let $x_{i}\in X_{i}$ for $i<n$, let $f_{i}=\frac{1}{2}$ for each $i$, and let $1_{n}=(1,1,\dots,1,0,0,\dots)\in\{0,1\}^{\mathbb{N}}$ be such that $(1_{n})_{i}=1$ exactly when $i<n$. For each $\beta\in\{0,1\}^{n}$, let $\beta^{\prime}\in\{0,1\}^{\mathbb{N}}$ be the sequence extending $\beta$ with zeroes, that is, $\beta^{\prime}_{i}=0$ for $i\geq n$. For one more piece of notation, we let
$c_{\beta^{\prime}}$ $=\frac{1}{8}\sum_{i=0}^{n}\beta^{\prime}_{i}\cdot 2^{-i}\textrm{ and }$
$c_{1_{n}}$ $=\frac{1}{8}\sum_{i=0}^{n}{(1_{n})}_{i}\cdot 2^{-i},$
noting that $0\leq c_{\beta^{\prime}},c_{1_{n}}<\frac{1}{4}$ are both rational.
@@
and if $\beta_{i}=1$, we have both $h_{\beta^{\prime}}(x_{i}),h_{1_{n}}(x_{i})\geq\frac{3}{4}$, so $h^{\prime}_{\beta}(x_{i})=\lambda\cdot h_{\beta^{\prime}}(x_{i})+(1-\lambda)\cdot h_{1}(x_{i})\geq\frac{3}{4}$. ∎
We now show that the logic perspective in some sense “fixes” these anomalies. If we consider “global learnability of a structure” – realizable PAC learnability for all definable families in a model – then there is no difference between the realizable and agnostic case, and we do have preservation under moving to statistical structures.
Proposition 6.
For any CL structure $\mathfrak{C}$, and definable predicate $\phi(x;y)$ with infinite $\gamma$ fat-shattering dimension for some $\gamma>0$, there is another definable predicate which defines a class that is not realizable PAC learnable. Thus in particular if every definable family for $\mathfrak{C}$ is realizable PAC learnable, then every definable family is agnostic PAC learnable, and thus every definable family in the randomized structure (and thus every distribution class or dual distribution class family) is agnostic PAC learnable and realizable PAC learnable.
We have stated the proposition for simplicity in the case where the partitioned formula has two free variables, but it extends to the general case of a partitioned formula.
Before beginning the proof of the proposition, we recall a basic result on the fat shattering dimension. If a class has infinite $\gamma$ fat-shattering dimension for some $\gamma>0$, this means we get arbitrary large powersets that we can capture with a $\gamma$-gap. The following result states that realize these counterexamples with a uniform choice of number $r<s$ that are at least $\gamma$-apart:
Fact 20 ((Alon et al., 1997, Thm. 4.1)).
Consider a real-valued hypothesis class ${\mathcal{H}}$ consisting of function $h_{y}$ for $y$ in some parameter set. Suppose that ${\mathcal{H}}$ has infinite $\gamma$ fat-shattering dimensions. Then for every natural number $d$, there are also $0\leq r<s\leq 1$ such that $s-r\geq\gamma$ and there are $x_{1},\dots,x_{d}\in X$ and $(y_{b}:b\in\{0,1\}^{d})$ such that for each $b$ and $i$, if $b(i)=0$, then $h(x_{i};y_{b})\leq r$, and if $b(i)=1$, then $h(x_{i};y_{b})\geq s$.
In the literature, this is sometimes phrased by saying that ${\mathcal{H}}$ has “infinite $V_{\gamma}$-dimension” (where $V$ is for Vapnik).
@@
Proof.
Fix $\mathfrak{C}$ and $\phi(x;y)$. Assume that $\phi(x;y)$ has infinite $\gamma$ fat-shattering dimension for some $\gamma>0$. Thus we get arbitrary large powersets that we can capture with a $\gamma$-gap. Then by Fact 20 for every natural number $d$, there are also $0\leq r<s\leq 1$ such that $s-r\geq\gamma$ and there are $x_{1},\dots,x_{d}\in X$ and $(y_{d}:b\in\{0,1\}^{d})$ such that for each $b$ and $i$, if $b(i)=0$, then $\phi(x_{i};y_{b})\leq r$, and if $b(i)=1$, then $\phi(x_{i};y_{b})\geq s$.
Let $f:[0,1]\to[0,1]$ be a continuous, monotone function such that $f(r)=0$ and $f(s)=1$. Then consider the real-valued function $\tau(x,y)$ defined as $f(\phi(x;y))$, which is also a definable predicate.
We now see that for every $d$, there are there are $x_{1},\dots,x_{d}$ in the domain of $\mathfrak{C}$ and $(y_{d}:b\in\{0,1\}^{d})$ such that for each $b$ and $i$, if $b(i)=0$, then $\tau(x_{i};y_{b})=0$, and if $b(i)=1$, then $\tau(x_{i};y_{b})=1$, so the $1$-graph dimension is at least $d$, so the class defined by $\tau(x;y)$ is not PAC learnable in the realizable case. ∎
@@ 6 Online learnability of statistical classes
We will now be interested in getting an analog of Theorem 5 for online learning, deriving bounds on learnability for statistical classes based on dimensions of the base class. We will calculate the following regret bound in terms of the sequential fat-shattering dimension:
Theorem 5.
For a function class ${\mathcal{H}}$ having $\frac{\epsilon}{50}$ sequential fat-shattering dimension at most $d$, the minimax regret of online learning for the randomization class of ${\mathcal{H}}$ over runs of length $n$ is at most
@@
If ${\mathcal{H}}$ is $\{0,1\}$-valued and has Littlestone dimension at most $d$, the minimax regret of online learning for the randomization class of ${\mathcal{H}}$ over runs of length $n$ is at most $O\left(\sqrt{d\cdot n}\right)$.
As finiteness of $\gamma$ sequential fat-shattering dimension for all $\gamma>0$ is equivalent to sublinear minimax regret by Fact 6, we have the following corollary:
Corollary 2.
If function class ${\mathcal{H}}$ is agnostic online learnable, in the sense of having sublinear minimax regret, then so is its randomization class, and thus so is its distribution class and the dual distribution class.
For the corollary, we use that the restrictions of the randomization class to deterministic range elements give the dual distribution class. We also use the fact that agnostic online learnability is closed under dualization by Proposition 1.
@@
To bound regret for online learning for the randomization class, we will adapt the framework of sequential Rademacher mean width developed in (Rakhlin et al., 2015a; b). We will define it here, slightly amending the definition to better match our conventions for mean width.
Definition 14 (Sequential Rademacher mean width).
@@
For a function $f(x,y)$, we then define the sequential Rademacher mean width, ${{\mathcal{R}}^{Seq}}_{f}(n)$, to be $\sup_{\bar{x}\in X^{\{1,-1\}^{<n}}}w_{{\mathcal{R}}^{Seq}}(f(\bar{x},Y))$. This definition coincides with the one from (Rakhlin et al., 2015a) except for the factor of $\frac{1}{n}$, although we call it a mean width instead of a complexity.
@@
Fact 21 (See (Rakhlin et al., 2015b, Proposition 9)).
@@
Theorem 6.
Let $(\Omega,\Sigma,\mu)$ be a probability space, and let $(f_{\omega}:\omega\in\Omega)$ be a family of real-valued functions $f:X\times Y\to[0,1]$ such that for each $(x,y)\in X\times Y$ and $v\in{\mathcal{R}}^{n}$, the functions $\omega\mapsto f_{\omega}(x,y)$ and $\omega\mapsto h_{f_{\omega}(\bar{x},Y)}(v)$ are measurable. Then
${{\mathcal{R}}^{Seq}}_{\mathbb{E}[f]}(n)\leq\sup_{\omega}{{\mathcal{R}}^{Seq}}_{f_{\omega}}(n).$
Proof.
Recall that the definition of sequential Rademacher mean width is the same as Rademacher mean width, except with a different probability distribution.
@@
$\sup_{\bar{x}\in X^{\{-1,1\}^{<n}}}w(\mathbb{E}_{\mu}[f_{\omega}](\bar{x},Y),\beta)\leq\sup_{\omega}\sup_{\bar{x}\in X^{\{-1,1\}^{<n}}}w(f_{\omega}(\bar{x},Y),\beta),$
and the proof of this is essentially identical to the proof of Theorem 1, but with a new distribution $\beta$. ∎
From this, we can prove a bound on regret for the randomization class in terms of a single sequential fat-shattering dimension, proving the first part of Theorem 5.
Theorem 7.
The minimax regret of online learning for the randomization class of ${\mathcal{H}}$ with $\gamma$ sequential fat-shattering dimension at most $d$ on a run of length $n$ is at most
@@
Proof.
By Theorem 6, any bound on ${{\mathcal{R}}^{Seq}}_{\mathcal{H}}(n)$ also applies to the randomization class, so regret for the randomization class is bounded by
@@
Because the function in the integral is decreasing in $\beta$, we may bound it naïvely by the value at $\gamma$. ∎
In case ${\mathcal{H}}$ is a $\{0,1\}$-valued concept class, all sequential fat-shattering dimensions coincide with the Littlestone dimension, and the following improved bound holds:
Fact 22 ((Alon et al., 2021, See Lemma 6.4 and Theorem 12.2)).
If $\mathcal{H}$ is a $\{0,1\}$-valued concept class with Littlestone dimension at most $d$, then
@@
From this, we are able to conclude that the minimax regret for the randomization class of $\mathcal{H}$ is also at most $O(\sqrt{d\cdot n})$, which proves the second part of Theorem 5.
6.2 Preservation of online learnability in moving to statistical classes, via logic
Above we showed that online learnability is preserved in moving to statistical classes. Without the quantitative bounds, this can also be derived from prior results in model theory. We can show that agnostic online learnability is equivalent to the notion of stability — a strengthening of NIP — for real-valued logic. We can then use prior results (Ben Yaacov, 2009; 2013) on the preservation of stability in real-valued logic under randomization. We now explain how to analyze online learnability via notions from logic, beginning with online learnable concept classes coming from first-order formulas over classical structures.
Consider a first-order structure $\mathfrak{M}$ and a formula with partitioned variables $\phi(\vec{x},\vec{y})$. We say $\phi$ is stable in $\mathfrak{M}$ if there do not exist arbitrarily large sequences $\vec{a}_{i},\vec{b}_{i}:1\leq n$ such that $\forall i,j\leq n$ $\phi(\vec{a}_{i},\vec{b}_{j})$ if and only if $i<j$. Roughly speaking stability says that $\phi$ does not define arbitrarily large linear orders.

### v2

@@ 4.5 Extension to the real-valued case
$\frac{\log^{2}(an)}{bn}\leq\frac{\left(\log C+\log(a^{2}b^{-1})+\log\log^{2}(ab^{-1})\right)^{2}}{bC(ab^{-1})\log^{2}(ab^{-1})}\leq\frac{\left(\log C+4\log(ab^{-1})\right)^{2}}{aC\log^{2}(ab^{-1})}$
$\leq\frac{(\log C/\log(ab^{-1})+4)^{2}}{C}\leq\frac{(\log C+4)^{2}}{C},$
@@
which completes the proof.
Theorem 15 for general real-valued hypothesis classes follows from Theorem 28 using Fact 10.
4.6 Simpler arguments for agnostic PAC learnability of the distribution function class
A much simpler argument is available that provides bounds for the distribution class or dual distribution class, but which does not extend to provide a bound for general measurable families as in Theorem 15. We explain the idea for the dual distribution class formed over a concept class. We know that if a concept class is agnostic PAC learnable, then so is its dual class, where the concepts are given by elements $x$ of the range space. We can then conclude by routine calculation with dimensions, that for each $k$, the functions on the parameter space given by normalized sums of elements $\frac{x_{1}+\ldots+x_{k}}{k}$ is agnostic PAC learnable. But by a basic result in statistical learning theory, the fact that our original concept class is PAC learnable means we can approximate each function in the dual distribution class arbitrarily closely by normalized sums.
Recall that in the special case of a concept class, the result for the dual distribution class had been proven in prior work (Hu et al., 2022). The simpler proof we present here, like the analytic proof that goes via the expectation class of a measurable family, improves on the bound given in prior work. We now explain the idea of this alternative approach.
@@ 4.6 Simpler arguments for agnostic PAC learnability of the distribution function class
This fits the composition framework in Theorem 1 of (Attias and Kontorovich, 2024).
Fact 16
Attias and Kontorovich (2024) The $\gamma$ fat-shattering dimension of the composed class $\chi^{\mathcal{C}}_{m}$ is bounded by:
@@
We now show that this approach generalizes easily from concept classes to function classes. That is, we give an alternative proof of the preservation of agnostic PAC learnability, without going through the expectation class of a measurable family. We do not compute sample complexity bounds explicitly for this alternative proof, but they are similar to those given above.
Theorem 29
@@
This again fits the composition framework in Theorem 1 of (Attias and Kontorovich, 2024). From the theorem we have the $\gamma$ fat-shattering dimension of the composed class is bounded by:
@@ 4.7 The realizable case for PAC learning
The previous subsections showed preservation of agnostic PAC learnability for the distribution class construction. We now show by contrast that the distribution class constructions do not preserve PAC learnability in the realizable case.
Proposition 30
There is a hypothesis class ${\mathcal{H}}$ that is realizable PAC learnable, but the distribution function class and dual distribution function class based on ${\mathcal{H}}$ are not realizable PAC learnable.
@@
Proof  A dual class element is given by an $x_{0}$ in the range space, with any other element in the same partition $X_{i}$ as $x_{0}$ inducing the same function. If we see samples $\vec{b}_{1}\ldots\vec{b}_{n}$, with value at most $\frac{1}{4}$, we will be able to exclude some partition elements $X_{i}$, but we will have infinitely many $X_{i}$ possible.
@@
Lemma 32
The distribution class of ${\mathcal{H}}_{0}$ is not PAC learnable in the realizable case. In fact, we may simply look at the class on $X$ of “two choice distributions”, consisting of all hypotheses $\lambda\cdot h_{b}+(1-\lambda)\cdot h_{b^{\prime}}$ for rational $\lambda\in[0,1]$ with $b,b^{\prime}\in B$.
Proof  We prove this by showing that this new class has infinite $\frac{1}{8}$-graph dimension, as defined in (Attias et al., 2023), which we now review. Fix a natural number $n$. For our purposes, we only need to know when a class has $\frac{1}{8}$-graph dimension at least $n$. The definition of graph dimension states that this holds when we have $x_{0},\dots,x_{n-1}\in X$, $f_{1},\dots,f_{n-1}\in[0,1]$, and for each $\beta\in\{0,1\}^{n}$, a hypothesis $h^{\prime}_{\beta}$ such that
@@
Note that, unlike with fat-shattering, we have an equality for the zero values of a branch and an inequality for the one value. Theorem 1 of (Attias et al., 2023) shows that an infinite $\frac{1}{8}$-graph dimension — that is, having such witnesses for each $n$ — implies that the class is not realizable PAC learnable. To find the required witnesses, let $x_{i}\in X_{i}$ for $i<n$, let $f_{i}=\frac{1}{2}$ for each $i$, and let $1_{n}=(1,1,\dots,1,0,0,\dots)\in\{0,1\}^{\mathbb{N}}$ be such that $(1_{n})_{i}=1$ exactly when $i<n$. For each $\beta\in\{0,1\}^{n}$, let $\beta^{\prime}\in\{0,1\}^{\mathbb{N}}$ be the sequence extending $\beta$ with zeros, that is, $\beta^{\prime}_{i}=0$ for $i\geq n$. For $b\in\{0,1\}^{\mathbb{N}}$, let
$c_{b}=\frac{1}{8}\sum_{i=0}^{\infty}b_{i}\cdot 2^{-i},$
so that for any $x_{i}\in X_{i},$
$h_{b}(x_{i})=\frac{3}{4}b_{i}+c_{b}.$
We note that for all $b$,
$0\leq c_{b}\leq\frac{1}{8}\sum_{i=0}^{\infty}2^{-i}=\frac{1}{4},$
and if there are only finitely many $i$ such that $b_{i}=1$, as is the case if $b=\beta^{\prime}$ for some $\beta\in\{0,1\}^{n}$ or $b=1_{n}$ for some $n$, then $c_{b}$ is rational.
@@
and if $\beta_{i}=1$, we have both $h_{\beta^{\prime}}(x_{i}),h_{1_{n}}(x_{i})\geq\frac{3}{4}$, so $h^{\prime}_{\beta}(x_{i})=\lambda\cdot h_{\beta^{\prime}}(x_{i})+(1-\lambda)\cdot h_{1}(x_{i})\geq\frac{3}{4}$. Thus in particular $|h^{\prime}_{\beta}(x_{i})-f_{i}|>\frac{1}{8}$ as required.
We now show that we can “fix” these anomalies by dealing not with a single hypothesis classes, but a family with reasonable closure properties. If we consider classes closed under composition with continuous bijections $[0,1]\to[0,1]$, then there is no difference between the realizable and agnostic case, and we do have preservation under moving to statistical classes.
Proposition 33
Let ${\mathcal{F}}$ be a family of hypothesis classes on range space $X$ parametrized by $Y$ closed under applying continuous bijections: for any ${\mathcal{H}}\in{\mathcal{F}}$, the composition of ${\mathcal{H}}$ with a continuous bijection $f$ is another class in ${\mathcal{F}}$. If ${\mathcal{F}}$ contains a hypothesis class with infinite $\gamma$ fat-shattering dimension for some $\gamma>0$, there is another hypothesis class in ${\mathcal{F}}$ which is not realizable PAC learnable. Thus in particular if every hypothesis class in ${\mathcal{F}}$ is realizable PAC learnable, then every class in ${\mathcal{F}}$ is agnostic PAC learnable, and thus every distribution class or dual distribution class arising from a class in ${\mathcal{F}}$ is agnostic PAC learnable and realizable PAC learnable.
Before beginning the proof of the proposition, we recall a basic result on the fat shattering dimension. If a class has infinite $\gamma$ fat-shattering dimension for some $\gamma>0$, this means we get arbitrary large powersets that we can capture with a $\gamma$-gap. The following result states that realize these counterexamples with a uniform choice of number $r<s$, where $s-r$ is bounded below by a function of $\gamma$:
Fact 18 (Alon et al. (1997, Thm. 4.2))
For every $\gamma>0$, there is some $\beta>0$ such that the following holds:
Consider a real-valued hypothesis class ${\mathcal{H}}$ consisting of functions $h_{y}$ for $y$ ranging over parameter set. Suppose that ${\mathcal{H}}$ has infinite $\gamma$ fat-shattering dimension. Then for every natural number $d$, there are also $0\leq r<s\leq 1$ such that $s-r\geq\beta$ and there are $x_{1},\dots,x_{d}\in X$ and $(y_{b}:b\in\{0,1\}^{d})$ such that for each $b$ and $i$, if $b(i)=0$, then $h(x_{i};y_{b})\leq r$, and if $b(i)=1$, then $h(x_{i};y_{b})\geq s$.
In the literature, this is sometimes phrased by saying that ${\mathcal{H}}$ has “infinite $V_{\beta}$-dimension” (where $V$ is for Vapnik).
@@
Proof  Fix a family ${\mathcal{F}}$ of hypothesis classes that is closed under applying continuous bijections. Assume there is a class ${\mathcal{H}}\in{\mathcal{F}}$ that is not agnostic PAC learnable, and we will prove that there is some ${\mathcal{H}}^{\prime}\in{\mathcal{F}}$ that is not realizable PAC learnable.
In particular, we assume that ${\mathcal{H}}$ has infinite $\gamma$ fat-shattering dimension for some $\gamma>0$. Then by Fact 18, there is some $\beta>0$ such that for every natural number $d$, there are also $0\leq r<s\leq 1$ such that $s-r\geq\beta$ and there are $x_{1},\dots,x_{d}\in X$ and $(y_{b}:b\in\{0,1\}^{d})$ such that for each $b$ and $i$, if $b(i)=0$, then ${\mathcal{H}}_{y_{b}}(x_{i})\leq r$, and if $b(i)=1$, then ${\mathcal{H}}_{y_{b}}(x_{i})\geq s$.
Let $f:[0,1]\to[0,1]$ be a continuous bijection such that $f(r)=0$ and $f(s)=1$. Then consider the hypothesis class =${\mathcal{H}}^{\prime}=\{h^{\prime}_{b}:b\in Y\}$ defined by $h^{\prime}_{b}(x)=f(h_{b}(x))$. Because ${\mathcal{F}}$ is closed under applying continuous bijections, ${\mathcal{H}}^{\prime}\in{\mathcal{F}}$, and we will show that ${\mathcal{H}}$ is not realizable PAC learnable.
To do this, we show that the $\frac{1}{8}$-graph dimension of ${\mathcal{H}}^{\prime}$ is infinite. (In fact, the $1-$graph dimension is.) For every $d$, there are there are $x_{1},\dots,x_{d}\in X$ and $(y_{b}:b\in\{0,1\}^{d})$ such that for each $b$ and $i$, if $b(i)=0$, then $h^{\prime}_{y_{b}}(x_{i})=0$, and if $b(i)=1$, then $h_{y_{b}}(x_{i})=1$.
To show that the $\frac{1}{8}$-graph dimension is greater than $d$, by the characterization in the proof of Lemma 32, it now suffices to define $f_{1},\dots,f_{n-1}\in[0,1]$ to all be 00, which ensures that
  • $h^{\prime}_{y_{b}}(x_{i})=f_{i}$ when $b_{i}=0$
  • $|h^{\prime}_{y_{b}}(x_{i})-f_{i}|>\frac{1}{8}$ when $b_{i}=1$,
because $|h^{\prime}_{y_{b}}(x_{i})-f_{i}|=h^{\prime}_{y_{b}}(x_{i})$.
@@ 5 Online learnability of statistical classes
We will now be interested in getting an analog of Theorem 15 for online learning, deriving bounds on learnability for statistical classes based on dimensions of the base class. As we did in the agnostic PAC case, we will work via a broader result on measurable families. Recall from Fact 6 that agnostic PAc learnability is controlled by the sequential fat-shattering dimension of a hypothesis class. We will calculate the following regret bound on the expectation class of a measurable family in terms of a bound on sequential fat-shattering dimension for the members in the family:
Theorem 34
For any $\epsilon,\delta>0$, if ${\mathcal{F}}$ is a measurable family of hypothesis classes such that each class ${\mathcal{H}}\in{\mathcal{F}}$ has $\frac{\epsilon}{50}$ sequential fat-shattering dimension at most $d$, the minimax regret of online learning for ${\mathbb{E}}{\mathcal{F}}$ over runs of length $n$ is at most
@@
If the classes in ${\mathcal{F}}$ are $\{0,1\}$-valued and have Littlestone dimension at most $d$, the minimax regret of online learning for ${\mathbb{E}}{\mathcal{F}}$ over runs of length $n$ is at most $O\left(\sqrt{d\cdot n}\right)$.
Note that we can again apply Proposition 12 to apply this to the distribution class of a fixed hypothesis class that has finite sequential fat-shattering dimensions. As finiteness of $\gamma$ sequential fat-shattering dimension for all $\gamma>0$ is equivalent to sublinear minimax regret by Fact 6, we have the following corollary:
Corollary 35
If function class ${\mathcal{H}}$ is agnostic online learnable, in the sense of having sublinear minimax regret, then so are the distribution class and the dual distribution class.
For the corollary, we use the fact that given any class ${\mathcal{H}}$ of finite $\gamma$ sequential fat-shattering dimension $d$, every element of the the parameter randomized family of ${\mathcal{H}}$ (see Definition 9) has $\gamma$ sequential fat-shattering dimension at most $d$. So its expectation class, ${\mathcal{H}}(\mathrm{CompatParamRV}({\mathcal{H}}))$, as in Definition 10, is agnostic online learnable. The distribution class then embeds into ${\mathcal{H}}(\mathrm{CompatParamRV}({\mathcal{H}}))$. To deal with the dual distribution class, we also use the fact that agnostic online learnability is closed under dualization by Proposition 3.
@@
To bound regret for online learning for the expectation class, we will adapt the framework of sequential Rademacher mean width developed in (Rakhlin et al., 2015a, b). We will define it here, slightly amending the definition to better match our conventions for mean width.
Definition 36 (Sequential Rademacher mean width)
@@
For a hypothesis class ${\mathcal{H}}$ on $X$, we then define the sequential Rademacher mean width, ${{\mathcal{R}}^{Seq}}_{f}(n)$, to be $\sup_{\bar{x}\in X^{\{1,-1\}^{<n}}}w_{{\mathcal{R}}^{Seq}}({\mathcal{H}}(\bar{x},Y)$. This definition coincides with the one from (Rakhlin et al., 2015a) except for the factor of $\frac{1}{n}$, although we call it a mean width instead of a complexity.
@@
Fact 19 (See (Rakhlin et al., 2015b, Proposition 9))
@@
Theorem 37
Let ${\mathcal{F}}=({\mathcal{H}}_{\omega}:\omega\in\Omega)$ be a measurable family of hypothesis classes on $X$. Then
${{\mathcal{R}}^{Seq}}_{\mathbb{E}[f]}(n)\leq\sup_{\omega}{{\mathcal{R}}^{Seq}}_{{\mathcal{H}}_{\omega}}(n).$
Proof  Recall that the definition of sequential Rademacher mean width is the same as Rademacher mean width, except with a different probability distribution.
@@
$\sup_{\bar{x}\in X^{\{-1,1\}^{<n}}}w({\mathbb{E}}{\mathcal{F}}(\bar{x},Y),\beta)\leq\sup_{\omega}\sup_{\bar{x}\in X^{\{-1,1\}^{<n}}}w({\mathcal{H}}_{\omega}(\bar{x},Y),\beta),$
and the proof of this is essentially identical to the proof of Theorem 22, but with a new distribution $\beta$.
From this, we can prove a bound on regret for the expectation class in terms of the sequential fat-shattering dimension, proving the first part of Theorem 34.
Theorem 38
Let ${\mathcal{F}}=({\mathcal{H}}_{\omega}:\omega\in\Omega)$ be a measurable family of hypothesis classes on $X$.
The minimax regret of online learning for ${\mathbb{E}}{\mathcal{F}}$ with $\gamma$ sequential fat-shattering dimension at most $d$ on a run of length $n$ is at most
@@
Proof  By Theorem 37, any uniform (in $\omega$) bound on ${{\mathcal{R}}^{Seq}}_{{\mathcal{H}}_{\omega}}(n)$ also applies to the expectation class, so regret for the expectation class is bounded by
@@
Because the function in the integral is decreasing in $\beta$, we may bound it naïvely by the value at $\gamma$.
For a $\{0,1\}$-valued concept class, all sequential fat-shattering dimensions coincide with the Littlestone dimension, and the following improved bound holds:
Fact 20 (Alon et al. (2021, See Lemma 6.4 and Theorem 12.2))
If ${\mathcal{C}}$ is a $\{0,1\}$-valued concept class with Littlestone dimension at most $d$, then
@@
From this, we are able to conclude that the minimax regret for the expectation class of a family of concept classes with Littlestone dimension bounded by $d$ is also at most $O(\sqrt{d\cdot n})$, which proves the second part of Theorem 34.
5.2 Preservation of online learnability in moving to statistical classes, via stability
Above we showed that online learnability is preserved in moving to statistical classes. We now give a variant without the quantitative bounds, but one which derives from prior results stated in the context of model theory. We can show that agnostic online learnability is equivalent to the notion of stability, a notion originating in logic. We can then use prior results (Ben Yaacov, 2009, 2013) on the preservation of stability in moving to an expectation class.
Definition 39
A concept class ${\mathcal{C}}$ over $X$ parameterized by $Y$ is stable if there do not exist arbitrarily large sequences $a_{i},b_{i}:1\leq n$ with $a_{i}\in X,b_{i}\in Y$ such that $\forall i,j\leq n$ $a_{i}\in{\mathcal{C}}_{b_{j}}$ if and only if $i<j$.
Roughly speaking stability says that ${\mathcal{C}}$ does not define arbitrarily large linear orders. This can be phrased as ${\mathcal{C}}$ having finite threshold dimension, as it is called in Ghazi et al. (2021), where this dimension is the largest size of a finite sequence of pairs $(a_{i},b_{i})$ with the above property.